OpenStudy (anonymous):

1+1/x=20/x^2 solve the equation

OpenStudy (saifoo.khan):

\[1+ \frac 1x = \frac{20}{x^2} \]Like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I kind of know how to do it but Im just lost in one part :/

OpenStudy (saifoo.khan):

Start with taking the lcm on left hand side of the equation.

OpenStudy (anonymous):

how do I get this \[X ^{2}=\left( -x ^{2}+20 \right)/x ^{2}\]

OpenStudy (saifoo.khan):